# Learning Heuristics for Transit Network Design and Improvement with Deep Reinforcement Learning

### v2

@@ 4.1.1 Training
Following the work of Kool et al. (2019), we train the policy net using REINFORCE with baseline, a policy gradient method proposed by Williams (1992), and set $\gamma=1$. Since, by construction, the reward $R_{t}=0$ except at $t_{\text{end}}$, this implies the return $G_{t}$ is the same for all steps:
$G_{t}=\sum^{t_{\text{end}}}_{t^{\prime}}R_{t^{\prime}}=-C(\mathcal{C},\mathcal{R})$ (9)
The learning signal for each action $a_{t}$ is $L_{t}=G_{t}-b(s_{t})$, where the baseline $b(s_{t})$ is a learned quantity that estimates the average $G_{t}$ achieved by $\pi_{\theta}$ starting in state $s_{t}$. Differing from Kool et al. (2019), we compute the baseline at each step using a small Multi-Layer Perceptron (MLP) neural net that is separate from the $\pi_{\theta}$. This baseline net is trained to predict $G_{t}$ based on the cost weight parameter $\alpha$ and statistics of the city $\mathcal{C}$.
We train the policy net on a dataset of synthetic cities. We sample a different $\alpha\sim[0,1]$ for each city, while holding $S,MIN,MAX$, and the constraint weight $\beta$ constant across training. The values of these parameters used are presented in Table 7. At each iteration, a full MDP episode is run on a “batch” of cities from the dataset, $C(\mathcal{C},\mathcal{R})$, $b(s_{t})$ and $L_{t}$ are computed across the batch, and back-propagation and parameter updates are applied to both the policy net (taking the gradient with respect to $-L_{t}\ln\pi_{\theta}$) and the baseline net (taking the gradient with respect to the mean squared error between $G_{t}$ and $b(s_{t})$).
@@
During training, we make a 90:10 split of this dataset into training and validation sets. Training progresses in “epochs”, where one epoch corresponds to running learned construction with the policy net on one MDP episode over each city in the training set. After each epoch, we run learned construction with the policy net on each city in the validation set, and the average cost of the resulting networks is recorded. At the end of training, the parameters $\theta$ from the epoch with the lowest average validation cost are returned, giving the final policy $\pi_{\theta}$.
@@ 4.2 Evolutionary Algorithm
Our baseline evolutionary algorithm is based on that of Nikolić and Teodorović (2013), with several modifications that we found improved its performance. We chose this algorithm because its ease of implementation and short running times (with appropriate parameters) aided prototyping and experimentation, but we note that our learned policies can in principle be used as heuristics in a wide variety of metaheuristic algorithms.
@@ 5.1 Comparison with Baseline Evolutionary Algorithm
Table 2 displays the mean and standard deviation of the cost $C(\mathcal{C},\mathcal{R})$ achieved by each algorithm on the Mandl and Mumford benchmarks for the operator perspective ($\alpha=0.0$), the passenger perspective ($\alpha=1.0$), and a balanced perspective ($\alpha=0.5$). We see that on Mandl, the smallest of the five cities, EA and NEA perform virtually identically for the operator and passenger perspectives, while NEA offers a slight improvement for the balanced perspective. But for the larger cities of the Mumford dataset, NEA performs considerably better than EA for both the operator and balanced perspectives. On Mumford3, the largest of the five cities, NEA solutions have 14% lower average cost for the operator perspective and 6% lower for the balanced perspective than EA.
@@
(a) Mumford1
(b) Mumford2
(c) Mumford3
Figure 5: Trade-offs between average trip time $C_{p}$ (on the x-axis) and total route time $C_{o}$ (on the y-axis) achieved by PC-EA, plotted along with LC-100, EA, and NEA (repeated from Figure 3) for comparison.
@@ 5.2.2 Contribution of type-2 mutator
We next considered the impact of the type-2 mutator heuristic on NEA. This low-level heuristic is the same between EA and NEA and is not a learned function. To understand how important it is to NEA’s performance, we ran “all-1 NEA”, a variant of NEA in which only the neural mutator is used, and not the type-2 mutator. The results are shown in Figure 4, along with the curve for NEA as in Figure 3, for comparison. It is clear that NEA and all-1 NEA perform very similarly in most scenarios. The differences are most notable at the extreme values of $\alpha$ (at either end of each curve), where we see that all-1 NEA underperforms NEA, achieving higher values of the relevant cost function component; but at intermediate $\alpha$ values on Mumford1 and Mumford2, NEA and all-1 NEA’s trip-time-versus-operating-cost curves overlap. It appears that the minor adjustments to existing routes made by the type-2 mutator are more important at extremes of $\alpha$, where the neural net policy has been pushed to the extremes of its behaviour.
On Mumford3, we observe that the curves do not overlap as cleanly; particularly for $\alpha\leq 0.5$, NEA seems to slightly outperform all-1 NEA. So the contribution of the type-2 mutator appears to grow as the city gets larger and more challenging. This may be because as the city gets larger, the space of possible routes grows larger, so the neural mutator’s changes get more dramatic in comparison with the single-node adjustments of the type-2 mutator. The type-2 mutator’s class of changes becomes more distinct from those of the neural mutator, possibly making it more important.
While the type-2 mutator’s contribution is small, it does help performance in some cases. We conclude that it is a useful component of the algorithm.
@@ 5.2.3 Importance of learned heuristics
We note that the type-1 mutator used in EA and the neural mutator used in NEA differ in the space of changes each is capable of making. The type-1 mutator can only add shortest paths as routes, while the neural mutator may compose multiple shortest paths into its new routes. It may be that this structural difference, as opposed to the quality of heuristics learned by the policy $\pi_{\theta}$, is part of NEA’s advantage over EA.
@@
We call this variant the path-combining evolutionary algorithm (PC-EA). Since we wanted to gauge the performance of this variant in isolation from the learned policy, we did not use LC-100 with $\pi_{\theta}$ to generate the initial network $\mathcal{R}_{0}$. Instead, we used $\pi_{\text{random}}$ in LC-100 to generate $\mathcal{R}_{0}$ for PC-EA.
The results are shown in Figure 5, along with the same LC-100, EA, and NEA curves as in Figure 3. For most values of $\alpha$, PC-EA performs less well than both EA and NEA. But at the extreme of $\alpha=1.0$ it performs better than both NEA and EA, decreasing average trip time $C_{p}$ by between one and two minutes versus NEA on each city - an improvement of about 5% on Mumford1 and Mumford3 and about 3.5% on Mumford2.
PC-EA’s poor performance at low $\alpha$ makes sense: its ability to replace routes with composites of shortest paths biases it towards making longer routes, moreso than EA, and unlike $\pi_{\theta}$, $\pi_{\text{random}}$ cannot learn to prefer halting early in such cases. This is an advantage over EA at $\alpha=1.0$ but a disadvantage at $\alpha$ less than about $0.8$. It is interesting, though, that PC-EA outperforms NEA at $\alpha=1.0$, as these two share the same space of possible actions. We would expect that the neural net policy used in NEA should be able to learn to reach that performance of PC-EA simply by learning to choose actions uniformly at random. Yet somehow it fails to do so.
Nonetheless, NEA performs better overall than PC-EA, and so we leave the exploration of this mystery to future work. For now we merely note that composing shortest paths is a good heuristic for route construction at high $\alpha$, but is worse than simply choosing shortest paths at low $\alpha$; so the advantage of NEA comes primarily from what the policy $\pi_{\theta}$ learns during training.
@@ 5.3 Comparisons with prior work
While conducting the experiments of subsection 5.1, we observed that after 400 iterations the cost of NEA’s best-so-far network $\mathcal{R}_{\text{best}}$ was still decreasing from one iteration to the next. In order to make a fairer comparison with other methods from the literature on the Mandl and Mumford benchmarks, we performed a further set of experiments where we ran NEA with $IT=4,000$ instead of $400$, for the operator perspective ($\alpha=0.0$) and the passenger perspective ($\alpha=1.0$). In addition, because we observed better performance from PC-EA for $\alpha$ near $1.0$, we also ran PC-EA with $IT=4,000$ for $\alpha=1.0$, and we include these results in the comparison.
@@
Table 3 and Table 4 present the results of these experiments on the Mandl and Mumford benchmarks, alongside results reported on these benchmarks in comparable recent work. As elsewhere, the results we report for NEA and PC-EA are averaged over ten runs with different random seeds; the same set of ten trained policies used in the experiments of subsection 5.1 and subsection 5.2 are used here in NEA. Results from other work are as reported in that work. In addition to the average trip time $C_{p}$ and total route time $C_{o}$, the tables present metrics of how many transfers between routes were required by the networks. These are labelled $d_{0}$, $d_{1}$, $d_{2}$, and $d_{un}$. $d_{i}$ with $i\in\{0,1,2\}$ is the percentage of all passenger trips that required $i$ transfers between routes; while $d_{un}$ is the percentage of trips that require more than 2 transfers ($d_{un}=100-(d_{0}+d_{1}+d_{2})$).
In these comparisons, we do not include results from Ahmed et al. (2019). While they reported results that set a new state of the art on these benchmarks, they do not provide the values used for the two parameters to their algorithm. In our earlier work Holliday and Dudek (2023), theirs was one of several methods we implemented in order to test our proposed initialization scheme. We discovered parameter values on our own that gave results comparable to other work up to 2019. But unfortunately, after much effort and correspondence with one of the authors of that work, we were unable to replicate their new state-of-the-art results. For this reason, we do not include their reported results here.
@@
Before discussing these results, a brief word should be said about computation time. On a desktop computer with a 2.4 GHz Intel i9-12900F processor and an NVIDIA RTX 3090 graphics processing unit (used to accelerate our neural net computations), NEA takes about 3 hours for each run on Mumford3, the largest environment, while the PC-EA runs take about 2.5 hours. By comparison, John et al. (2014) and Hüsselmann et al. (2023) both use a variant of NSGA-II, a genetic algorithm, with a population of 200 networks, which in both cases takes more than two days to run on Mumford3. Hüsselmann et al. (2023)’s DBMOSA variant, meanwhile, takes 7 hours and 52 minutes to run on Mumford3. Kılıç and Gök (2014) report that their procedure takes eight hours just to construct the initial network for Mumford3, and don’t report the running time for the subsequent optimization.
@@
These reported running times are not directly comparable as the experiments were not run on identical hardware, but the differences are broadly indicative of the differences in speed of these methods. The running times of these methods is mainly a function of the metaheuristic algorithm used, rather than the low-level heuristics used in the algorithm. Genetic algorithms like NSGA-II, with large populations as used in some of these methods, are very time-consuming because of the large number of networks they must modify and evaluate at each step. But their search of solution space is correspondingly more exhaustive than single-solution methods such as simulated annealing, or an evolutionary algorithm with a small population ($B=10$) as we use in our own experiments. It is therefore to be expected that they would achieve lower final costs in exchange for their greater run-time. We are interested here in the quality of the low-level heuristics we propose, rather than the metaheuristic algorithm, and this must be kept in mind as we discuss the results of this section.
City Method $C_{p}\downarrow$ $C_{o}\downarrow$ $d_{0}\uparrow$ $d_{1}$ $d_{2}$ $d_{un}\downarrow$
@@
NEA 29.27 5155 30.36 51.01 17.43 1.2
PC-EA 28.09 6830 38.6 57.02 4.35 0.03
Table 3: Passenger-perspective results. $C_{p}$ is the average passenger trip time. $C_{o}$ is the total route time. $d_{i}$ is the percentage of trips satisfied with number of transfers $i$, while $d_{un}$ is the percentage of trips satsified with 3 or more transfers. Arrows next to each quantity indicate which of increase or decrease is desirable. Bolded values in $C_{p}$, $d_{0}$, and $d_{un}$ columns are the best on that environment.
@@ 5.3.1 Passenger-perspective results
Table 3 shows the passenger perspective results alongside results from other work. On each city except Mumford1, the NSGA-II variant of Hüsselmann et al. (2023) has the best final performance; it is also among the most time-consuming of the algorithms reported on here. PC-EA and NEA both perform poorly on the smallest two cities, Mandl and Mumford0, but their relative performance improves as the size of the city increases. On both Mumford2 and Mumford3, PC-EA’s performance is very close to that of Hüsselmann et al. (2023)’s NSGA-II, and better than all other methods listed, despite the relatively under-powered metaheuristic that drives it.
PC-EA also outperforms Hüsselmann et al. (2023)’s DBMOSA, which uses the same low-level heuristics as their NSGA-II with a faster but less-exhaustive metaheuristic. That PC-EA exceeds DBMOSA’s performance on Mumford2 and Mumford3 is evidence that the low-level heuristics used in PC-EA are better for the passenger perspective than those proposed by Hüsselmann et al. (2023). This is especially the case given that DBMOSA is still a more sophisticated metaheuristic algorithm than ours, with ten different low-level heuristics and a hyper-heuristic that adapts the rate at which each low-level heuristic is applied over the run.
NEA does not perform as well as PC-EA on the passenger perspective (aligning with what we observed in subsection 5.2), but still shows good performance on Mumford2 and 3, outperforming all of these methods that were published prior to 2022.

### v3

@@ 4.1.1 Training
In our prior work (Holliday and Dudek, 2023, 2024), we trained the policy net using the REINFORCE with Baseline method proposed by Williams (1992), using a reward function that was 0 at all steps except the final step, where it was the negative of the cost function. This was inspired by the similar approach taken by Kool et al. (2019). However, REINFORCE has been been built upon by a number of other policy gradient learning algorithms since its publication. Also, sparse reward functions like this one, which is zero at most timesteps, convey little information about the effect of each action, and so can make learning difficult.
For these reasons, in this work we train the policy net using PPO, a policy gradient method proposed by Schulman et al. (2017), and use the richer reward function described in section 3.2. The PPO method was chosen for its relative simplicity and for the fact that it remains a state-of-the-art method on many problems. PPO works by rolling out a number of steps $H$ of a “batch” of episodes in parallel, and then computing an “advantage” for each timestep, $A_{t}=\sum^{H-1}_{t^{\prime}=0}\gamma^{t^{\prime}}R_{t+t^{\prime}}+\gamma^{H}V(s_{t+H})-V(s_{t})$. This advantage weights the updates to the policy, such that actions with a positive advantage are made more likely, and those with a negative advantage are made less so.
The value function $V(s_{t})$ is itself a small Multi-Layer Perceptron (MLP) neural net that is trained in parallel with the policy to predict the discounted return $G^{H}_{t}=\sum^{H-1}_{t^{\prime}=0}\gamma^{t^{\prime}}R_{t+t^{\prime}}+\gamma^{H}V(s_{t+H})$, based on the cost weight parameter $\alpha$ and statistics of the city $\mathcal{C}$ and of the current network $\mathcal{R}_{t},r_{t}$.
We train the policy net on a dataset of synthetic cities. We sample a different $\alpha\sim[0,1]$ for each city, while holding $S,MIN,MAX$, and the constraint weight $\beta$ constant across training. The values of these parameters used are presented in Table 7. At each iteration, a full MDP episode is run on a “batch” of cities from the dataset, $C(\mathcal{C},\mathcal{R})$, $V(s_{t})$ and $G^{H}_{t}$ are computed across the batch, and back-propagation and parameter updates are applied to both the policy net and the value net according to PPO.
@@
During training, we make a 90:10 split of this dataset into training and validation sets. In each iteration of training, we run learned construction with the policy net on one MDP episode over each city in a batch that is sampled from the dataset, and update the policy and value nets. After every ten iterations, we run learned construction with the policy net on each city in the validation set, and the average cost of the resulting networks is recorded. At the end of training, the parameters $\theta$ that achieved the lowest average validation cost are returned, giving the final policy $\pi_{\theta}$. Training proceeds for 200 iterations in batches of 256 cities, as we found that the policy stopped improving after this. A total of 51,200 cities are used during training: less than the entire training dataset.
@@ 4.2 Evolutionary Algorithm
Our baseline evolutionary algorithm is based on that of Nikolić and Teodorović (2013), with several modifications that we found improved its performance. We chose this algorithm because its ease of implementation and its good performance at short running times (with appropriate parameters), aided prototyping and experimentation, but we note that our learned policies can in principle be used as heuristics in a wide variety of metaheuristic algorithms.
@@ 5.1 Comparison with Baseline Evolutionary Algorithm
Table 2 displays the mean and standard deviation of the cost $C(\mathcal{C},\mathcal{R})$ achieved by each algorithm on the Mandl and Mumford benchmarks for the operator perspective ($\alpha=0.0$), the passenger perspective ($\alpha=1.0$), and a balanced perspective ($\alpha=0.5$). We see that on Mandl, the smallest of the five cities, EA and NEA perform virtually identically for the operator and passenger perspectives, while NEA offers a slight improvement for the balanced perspective. But for the larger cities of the Mumford dataset, NEA performs considerably better than EA for both the operator and balanced perspectives. On Mumford3, the largest of the five cities, NEA solutions have 13% lower average cost for the operator perspective and 5% lower for the balanced perspective than EA.
@@
Figure 5: Trade-offs between average trip time $C_{p}$ (on the x-axis) and total route time $C_{o}$ (on the y-axis) achieved by RC-EA, plotted along with NEA (repeated from Figure 3) for comparison.
@@ 5.2.2 Contribution of type-2 mutator
We next considered the impact of the type-2 mutator heuristic on NEA. This low-level heuristic is the same between EA and NEA and is not a learned function. To understand how important it is to NEA’s performance, we ran “all-1 NEA”, a variant of NEA in which only the neural mutator is used, and not the type-2 mutator. The results are shown in Figure 4, along with the curve for NEA as in Figure 3, for comparison. It is clear that NEA and all-1 NEA perform very similarly in most scenarios. The most notable difference is at $\alpha=1.0$, where we see that all-1 NEA underperforms NEA, achieving higher values of $C_{p}$. It appears that the minor adjustments to existing routes made by the type-2 mutator are more important at extremes of $\alpha$, where the neural net policy has been pushed to the extremes of its behaviour.
On Mumford3, we observe that the curves do not overlap as cleanly: NEA seems to slightly outperform all-1 NEA at most $\alpha$ values. So the contribution of the type-2 mutator appears to grow as the city gets larger and more challenging. This may be because as the city gets larger, the space of possible routes grows larger, so the neural mutator’s changes get more dramatic in comparison with the single-node adjustments of the type-2 mutator. The type-2 mutator’s class of changes becomes more distinct from those of the neural mutator, possibly making it more important.
While the type-2 mutator’s contribution is small, it does help performance in some cases, so we conclude that it is a useful component of the algorithm.
@@ 5.2.3 Importance of learned heuristics
We note that the type-1 mutator used in EA and the neural mutator used in NEA differ in the space of changes each is capable of making. The type-1 mutator can only add shortest paths as routes, while the neural mutator may compose multiple shortest paths into its new routes. It may be that this structural difference, as opposed to the quality of the heuristic learned by the policy $\pi_{\theta}$, is part of NEA’s advantage over EA.
@@
We call this variant the randomly-combining evolutionary algorithm (RC-EA). Since we wanted to gauge the performance of this variant in isolation from the learned policy, we did not use LC-100 with $\pi_{\theta}$ to generate the initial network $\mathcal{R}_{0}$. Instead, we used $\pi_{\text{random}}$ in LC-100 to generate $\mathcal{R}_{0}$ for RC-EA.
The results are shown in Figure 5, along with the same NEA curves as in Figure 3. For most values of $\alpha$, RC-EA performs less well than both EA and NEA. But at the extreme of $\alpha=1.0$ it performs better than NEA, decreasing average trip time $C_{p}$ by between one and two minutes versus NEA on each city - an improvement of about 3.7% on Mumford1 and about 4.5% on Mumford2 and Mumford3.
RC-EA’s poor performance at low $\alpha$ makes sense: its ability to replace routes with composites of shortest paths biases it towards making longer routes, and unlike $\pi_{\theta}$, $\pi_{\text{random}}$ cannot learn to prefer halting early in such cases. This is an advantage over NEA at $\alpha=1.0$ but a disadvantage at $\alpha$ less than about $0.8$. It is interesting, though, that RC-EA outperforms NEA at $\alpha=1.0$, as these two share the same space of possible actions. We would expect that the neural net policy used in NEA should be able to learn to reach that performance of RC-EA simply by learning to choose actions uniformly at random. Yet somehow it fails to do so.
Nonetheless, NEA performs better overall than RC-EA, and so we leave the exploration of this mystery to future work. For now we merely note that composing shortest paths is a good heuristic for route construction at high $\alpha$, but is worse than simply choosing shortest paths at low $\alpha$; so the advantage of NEA comes primarily from what the policy $\pi_{\theta}$ learns during training.
@@ 5.3 Comparisons with prior work
While conducting the experiments of subsection 5.1, we observed that after 400 iterations the cost of NEA’s best-so-far network $\mathcal{R}_{\text{best}}$ was still decreasing from one iteration to the next. In order to make a fairer comparison with other methods from the literature on the Mandl and Mumford benchmarks, we performed a further set of experiments where we ran NEA with $IT=4,000$ instead of $400$, for the operator perspective ($\alpha=0.0$) and the passenger perspective ($\alpha=1.0$). In addition, because we observed better performance from RC-EA for $\alpha$ near $1.0$, we also ran RC-EA with $IT=4,000$ for $\alpha=1.0$, and we include these results in the comparison.
@@
Table 3 and Table 4 present the results of these experiments on the Mandl and Mumford benchmarks, alongside results reported on these benchmarks in comparable recent work. As elsewhere, the results we report for NEA and RC-EA are averaged over ten runs with different random seeds; the same set of ten trained policies used in the experiments of subsection 5.1 and subsection 5.2 are used here in NEA. Results from other work are as reported in that work. In addition to the average trip time $C_{p}$ and total route time $C_{o}$, the tables present metrics of how many transfers between routes were required by the networks. These are labelled $d_{0}$, $d_{1}$, $d_{2}$, and $d_{un}$. $d_{i}$ with $i\in\{0,1,2\}$ is the percentage of all passenger trips that required $i$ transfers between routes; while $d_{un}$ is the percentage of trips that require more than 2 transfers ($d_{un}=100-(d_{0}+d_{1}+d_{2})$).
@@
Before discussing these results, a brief word should be said about computation time. On a desktop computer with a 2.4 GHz Intel i9-12900F processor and an NVIDIA RTX 3090 graphics processing unit (used to accelerate our neural net computations), NEA takes about 10 hours for each 4,000-iteration run on Mumford3, the largest environment, while the RC-EA runs take about 6 hours. By comparison, John et al. (2014) and Hüsselmann et al. (2023) both use a variant of NSGA-II, a genetic algorithm, with a population of 200 networks, which in both cases takes more than two days to run on Mumford3. Hüsselmann et al. (2023)’s DBMOSA variant, meanwhile, takes 7 hours and 52 minutes to run on Mumford3. Kılıç and Gök (2014) report that their procedure takes eight hours just to construct the initial network for Mumford3, and don’t report the running time for the subsequent optimization.
@@
These reported running times are not directly comparable as the experiments were not run on identical hardware, but they are broadly indicative of the differences in speed of these methods. The running times are mainly a function of the metaheuristic algorithm used, rather than the low-level heuristics used in the algorithm. Genetic algorithms like NSGA-II, with large populations as used in some of these methods, are very time-consuming because of the large number of networks they must modify and evaluate at each step. But their search of solution space is correspondingly more exhaustive than single-solution methods such as simulated annealing, or an evolutionary algorithm with a small population ($B=10$) as we use in our own experiments. It is therefore to be expected that they would achieve lower final costs in exchange for their greater run-time. We are interested here in the quality of our low-level heuristics, rather than that of the metaheuristic algorithm, and this must be kept in mind as we discuss the results of this section.
5.3.1 Caveat for Ahmed et al. (2019)
In these comparisons, we include results from Ahmed et al. (2019), but we wish to state a caveat to these. While they reported results that set a new state of the art on these benchmarks, they do not provide the values used for the two parameters to their algorithm. In our earlier work Holliday and Dudek (2023), theirs was one of several methods we implemented in order to test our proposed initialization scheme. We discovered parameter values on our own that gave results comparable to other work up to 2019, but after much effort and correspondence with one of the authors of Ahmed et al. (2019), we were unable to replicate their state-of-the-art results.
City Method $C_{p}\downarrow$ $C_{o}$ $d_{0}\uparrow$ $d_{1}$ $d_{2}$ $d_{un}\downarrow$
@@
NEA 29.18 5122 30.70 52.27 16.17 0.87
RC-EA 28.09 6830 38.6 57.02 4.35 0.03
Table 3: Passenger-perspective results. $C_{p}$ is the average passenger trip time. $C_{o}$ is the total route time. $d_{i}$ is the percentage of trips satisfied with number of transfers $i$, while $d_{un}$ is the percentage of trips satisfied with 3 or more transfers. Arrows next to each quantity indicate which of increase or decrease is desirable. Bolded values in $C_{p}$, $d_{0}$, and $d_{un}$ columns are the best on that environment.
@@ 5.3.2 Passenger-perspective results
Table 3 shows the passenger perspective results alongside results from other work. On each city except Mumford3, the hyper-heuristic method of Ahmed et al. (2019) performs best, while on Mumford3 the NSGA-II variant of Hüsselmann et al. (2023) performs best. RC-EA and NEA both perform poorly on the smallest two cities, Mandl and Mumford0, but their relative performance improves as the size of the city increases. On both Mumford2 and Mumford3, RC-EA’s performance is very close to that of Hüsselmann et al. (2023)’s and Ahmed et al. (2019)’s, and better than all other methods listed, despite the relatively under-powered metaheuristic that drives it.
RC-EA also outperforms Hüsselmann et al. (2023)’s DBMOSA, which uses the same low-level heuristics as their NSGA-II with a faster but less-exhaustive metaheuristic. That RC-EA exceeds DBMOSA’s performance on Mumford2 and Mumford3 is evidence that the low-level heuristics used in RC-EA are better for the passenger perspective than those proposed by Hüsselmann et al. (2023). This is especially the case given that DBMOSA is still a more sophisticated metaheuristic algorithm than ours, with ten different low-level heuristics and a hyper-heuristic that adapts the rate at which each low-level heuristic is applied over the run - though we acknowledge that DBMOSA is a multi-objective optimization algorithm, which may account for its relative weakness here.
NEA does not perform as well as RC-EA on the passenger perspective (aligning with what we observed in subsection 5.2), but still shows good performance on Mumford2 and 3, outperforming all of these methods that were published prior to 2019.